# PuzzleBench: Can LLMs Solve Challenging First-Order Combinatorial Reasoning Problems?

### v2

@@ 8 Qualitative Analysis
1) Problems that our solved with 100% accuracy without any feedback from solved examples. 8 such problems exist out of the 31, including vertex-cover and latin-square. These problems have a one-to-one correspondence between the natural language description of the rules and the program for generating the constraints and the LLM essentially has to perform a pure translation task.
@@
2) Problems that our solved with 100% accuracy but which need feedback from solved examples. There are 17 such problems. They typically do not have a one-to-one correspondence between rule descriptions and code, thus requiring some reasoning to encode the problem in the solver’s language. GPT-4-Turbo initially misses constraints or encodes the problem incorrectly, but with feedback, it can spot its mistakes and corrects its programs. Examples include k-clique and binairo. In binairo, for example, GPT-4-Turbo incorrectly encodes the constraints for ensuring all columns and rows are distinct. In k-clique, it interprets the problem incorrectly and adds an extra constraint to ensure that if a vertex is part of a clique, then all its neighbours are also a part of the clique.
@@
3) Problems with performance below 100% which are not corrected through feedback or utilizing multiple runs. For these 6 problems, LLM finds it difficult to encode some natural language constraint into SMT. Examples include number-link and hamiltonian path, where GPT-4-Turbo is not able to figure out how to encode existence of paths as SMT constraints. In our opinion, these conversions are peculiar, and may be hard even for average computer science students.
9 Conclusion and Future Work
@@
We explore the reasoning abilities of LLMs on challenging first-order combinatorial reasoning problems. We present PuzzleBench, a novel benchmark of 31 such problems and find that existing tool-augmented techniques fair poorly. In response, we present Puzzle-LM a new technique to aid LLMs with both program interpreters and symbolic solvers. Our extensive experiments show Puzzle-LM’s superior performance on our dataset. We also find that solved examples provide useful feedback to improve LLM reasoning. Further analyis reveals that Puzzle-LM struggles for a certain class of problems where conversion to symbolic representation is not straightforward. In such cases simple feedback strategies do not improve reasoning; exploring methods to alleviate such problems is a promising direction for future work. We release our code and data for further research.
10 Limitations
Dataset contamination is a big problem in the era of pretrained language models which have been trained on large web corpora. Therefore, its really hard to determine if a dataset has been seen as part of the training data. Determining the extent of contamination is also not easy. Since PuzzleBench does contain some standard problems from Computer Science literature, there is possibility of contamination. Usage of LLMs for reasoning on text is also limited by availability of large compute resources.
Our approach is also limited by the nature of the symbolic solver used. Some problems cannot be encoded in the symbolic language used by the solver, although our approach remains flexible enough to include several such solvers. Even in the case that a problem can be encoded in the symbolic language used by the solver, this encoding can potentially be complicated. Our method relies on the in-context learning ability of LLMs to achieve this. Additionally we assume that input instances and their outputs have a fixed pre-defined (serialized) representation.
11 Ethics Statement
Strubell et al. (2019) show that the financial and environmental costs that are associated with the training of LLMs lead to significant contribution to global warming. We don’t train the model from scratch and leverage in context learning, so the energy footprint of our work is less. The large language models whose API we use for inference, especially GPT-4, consume significant energy.
Acknowledgements
Mausam is supported by grants from Google, Huawei, and Jai Gupta Chair Professorship from IIT Delhi. Parag was supported by the DARPA Explainable Artificial Intelligence (XAI) Program (#N66001-17-2-4032). Mausam and Parag are also supported by IBM SUR awards. Any opinions, findings, conclusions, or recommendations expressed here are those of the authors and do not necessarily reflect the views or official policies, either expressed or implied, of the funding agencies. We also acknowledge the Microsoft Accelerate Foundation Models Research Program for GPT4 access.
Appendix A PuzzleBench
@@ A.1 Dataset Details and Statistics
Our dataset namely PuzzleBench has $31$ different fcore problems that have been collected from various sources. Some of these problems have been collected from Nikoli and other problems are well-known computational problems from Computer Science literature such as hamiltonian path and minimum-dominating set. Table 5 gives the details of all problems in our dataset. To create our training and test sets, we synthetically generate problem instances when possible and we handcraft problem instances otherwise. Each problem has some solved training instances and a separate set of testing instances. Each problem also has a natural language description of its rules, and a natural language description of the input-format which specify how input problem instances and their solutions are represented in text. The next few sections give illustrative examples and other details.
@@ A.2 Natural Language Description of Rules
This section describes how we create the natural language description of rules for problems in PuzzleBench. We extract rules from Wikipedia/Nikoli pages of the corresponding problems. These rules are reworded by a human expert to reduce dataset contamination. Another human expert ensures that there are no ambiguities in the reworded description of the rules. The rules are generalized, when needed (for eg. from a $9\times 9$ Sudoku to a $n\times n$ Sudoku). The following sections provide few examples.
@@ A.2.1 Example Problem: Survo
Survo (Figure 6) is an example problem from PuzzleBench. The task is to fill a $m\times n$ rectangular board with numbers from $1-m*n$ such that each row and column sums to an intended target. (Survo-Wikipedia). The box given below describes the rules of Survo more formally in natural language. We are given a partially filled $m\times n$ rectangular board, intended row sums and column sums. - Empty cells are to be filled with numbers - Numbers in the solved board can range from $1$ to $m*n$ Numbers present in filled cells on the input board cannot be removed - Each number from 1 to m*n must appear exactly once on the solved board - All the empty cells should be filled such that each row and each column of the solved board must sum to the respective row sum and column sum as specified in the input
@@ A.2.2 Example Problem: Hamiltonian Path
We are given an un-directed and un-weighted graph. - We have to determine if the graph contains a path that visits every vertex exactly once.
@@ A.2.3 Example Problem: Dosun Fuwari
Dosun Fuwari (Nikoli) as shown in Figure 8 is another example problem from PuzzleBench. We are given a square board with regions (cells enclosed in bold lines) and we have to fill the board with balloons and iron balls such that one balloon and one iron ball is placed in each region. Balloons are light and float, so they must be placed in one of the cells at the top, in a cell right under a black cell (filled-in cell), or under other balloons. Iron balls are heavy and sink, so they must be placed in one of the cells at the bottom, or in a cell right over a black cell or over other iron balls. The box given below gives the more formal description of the rules of dosun fuwari in natural language.
We are given a partially filled n*n square board. We are also given subgrids of the input board. Cells in the input board can either be empty or filled (that is, nothing can be placed in them, they are blackened) or can be balloons or iron balls. - The only thing we can do is place balloons or iron balls in some of or all of the empty cells - Each subgrid specified in the input should have exactly one balloon and iron ball in the solved board - Because balloons are buoyant, they should be positioned either in one of the cells located at the top of the board or in a cell directly below a filled cell (i.e., one of the blackened cells in the input) or below other balloons. - Iron balls, being dense, will sink and should therefore be positioned either directly on one of the cells located at the bottom of the input board, or on a cell directly above a filled cell (i.e., one of the blackened cells in the input), or above another iron ball.
@@ A.3 Natural Language Description of Input and Output format
For many problems we consider input-output instances are typically not represented in text. For each problem we describe a straightforward conversion of the input and output space to text in natural language. The following sections consider examples of a few problems from PuzzleBench.
@@ A.3.1 Example Problem: Survo
Input Format: - The input will have $m+1$ lines - The first $m$ lines will have $n+1$ space-separated integers - Each of these $m$ lines represents one row of the partially solved input board (n integers), followed by the required row sum (a single integer) - The last line of the input will have $n$ space-separated integers each of which represents the required column sum in the solved board Sample Input: 0 6 0 0 0 30 8 1 0 0 0 17 0 9 3 0 30 27 16 10 25 Output Format: - The output should have $m$ lines, each representing one row of the solved board - Each of these $m$ lines should have $n$ space-separated integers representing the cells of the solved board - Each integer should be from $1$ to $m*n$ Sample Output: 12 6 2 10 8 1 5 4 7 9 3 11
@@ A.3.2 Example Problem: Dosun Fuwari
Figure 10 represents conversion of the inputs to dosun fuwari, originally represented as grid images to text. Here the first few lines represent the input board followed by a string ’—–’ which acts as a separator following which each of the lines has space-separated integers representing the subgrids of the input board. Cells are numbered in row-major order starting from 0, and this numbering is used to represent cells in each of the lines describing the subgrids. In the first few lines representing the input board, 0’s represent the empty cells that must be filled. 1’s denote the blackened cell, 2s denote the balloons and 3’s denote the iron balls. The box below describes these rules more formally in natural language Input-Format: - The first few lines represent the input board, followed by a line containing --------, which acts as a separator, followed by several lines where each line represents one subgrid - Each of the lines representing the input board will have space-separated integers ranging from 0 to 3 - 0 denotes empty cells, 1 denotes a filled cell (blackened cell), 2 denotes a cell with a balloon, 3 denotes a cell with an iron ball - After the board, there is a separator line containing -------- - Each of the following lines has space-separated elements representing the subgrids on the input board - Each of these lines has integers representing cells of a subgrid - Cells are numbered in row-major order starting from 0, and this numbering is used to represent cells in each of the lines describing the subgrids Sample-Input: 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 1 -------- 0 1 2 3 6 10 4 8 12 5 9 13 14 15 7 11 Output Format: - The output should contain as many lines as the size of the input board, each representing one row of the solved board - Each row should have n space separate integers (ranging from 0-3) where n is the size of the input board - Empty cells will be denoted by 0s, filled cells (blackened) by 1s, balloons by 2s and iron balls by 3s Sample-Output: 2 3 0 2 2 1 0 2 0 2 3 3 3 0 3 1
@@ A.3.3 Example Problem: Hamiltonian Path
Figure 11 represents the conversion of inputs to hamiltonian-path, originally represented as graph image to text. The first line denotes the number of vertices present in the graph followed by which each node of the graph will be numbered from 0 - N-1. Each of the subsequent lines represents an edge of the graph and will contain two space-separated integers (according to the numbering defined previously). The output is a single word (YES/NO) indicating if a hamiltonian path exists in the graph. The box below describes this more formally in natural language. Input Format: - The first line will contain a single integer N, the number of nodes in the graph - The nodes of the graph will be numbered from 0 to N-1 - Each of the subsequent lines will represent an edge of the graph and will contain two space-separated integers (according to the numbering defined above) Sample-Input: 5 0 1 1 2 2 3 3 4 Output Format: - The output should contain a single line with a single word - The word should be YES if a path exists in the input graph according to constraints specified above and NO otherwise Sample Output: YES
@@
3-Partition (Non Decision)
15
30
12 (array size)
17.7
NP-Hard
3-Partition (Decision)
15
30
12 (array size)
17.7
NP-Complete
Binario
15
50
4.0$\times$4.0 (grid size)
6.96$\times$6.96
NP-Hard De Biasi (2013)
Clique Cover
15
30
6.26, 9.4 (# of nodes, # of edges)
12.9, 31.4
NP-Complete
Cryptarithmetic
15
25
4.32 (Average # of digits in the two operands )
3.26
NP-Hard Epstein (1987)
Dosun Fuwari
15
30
3.066$\times$3.066 (grid size)
5.23$\times$5.23
NP-Hard Iwamoto and Ibusuki (2018)
Futoshiki
15
47
5$\times$5 (grid size)
7.57$\times$7.57
NP-Hard Lloyd et al. (2022)
Graph Coloring
15
30
5.13, 6.8 (# of nodes, # of edges)
9, 21.06
NP-Complete Gent et al. (2017)
Hamiltonian Path
15
30
5.93, 8.6 (# of nodes, # of edges)
13.0, 19.77
NP-Complete
Hamiltonian Cycle
15
27
5.93, 8.6 (# of nodes, # of edges)
11.07, 18.67
NP-Complete
Independent Set
12
30
5.83, 7.25 (# of nodes, # of edges)
14.26, 29.8
NP-Complete
Inshi-No-Heya
15
49
5.0$\times$5.0 (grid size)
6.5$\times$6.5
K-Clique
15
31
4.87, 7.67 (# of nodes, # of edges)
8.84, 26.97
NP-Complete
Keisuke
15
30
4.33$\times$4.33 (grid size)
5.83$\times$5.83
Ken Ken
15
20
3.26$\times$3.26 (grid size)
5.2$\times$5.2
NP-Hard Haraguchi and Ono (2015)
Latin Square
15
50
6.0$\times$6.0 (grid size)
14.3$\times$14.3
NP-Hard Colbourn (1984)
Longest Path Problem
15
30
6.2, 5.87 (# of nodes, # of edges)
12.6, 16.3
NP-Complete
Magic Square
15
30
3.0$\times$3.0 (grid size)
4.33$\times$4.33
Minimum Dominating Set
15
30
6.0, 17.73 (# of nodes, # of edges)
14.53, 45.0
NP-Complete
Number Link
15
50
4$\times$4 (grid size)
7.1$\times$7.1
NP-Hard
Partition Problem
15
35
7.06 (array size)
15
NP-Complete
Shinro
15
25
5.13$\times$5.13 (grid size)
9.2$\times$9.2
Subset Sum
15
30
3.67 (array size)
11.87
NP-Complete
Sudoku
15
50
4.0$\times$4.0 (grid size)
13.3$\times$13.3
NP-Hard YATO and SETA (2003)
Sujiko
15
45
3.0$\times$3.0 (grid size)
4.0$\times$4.0
Survo
15
47
13.5 (area of grid)
20.25
Symmetric Sudoku
15
30
4$\times$4 (grid size)
6.5$\times$6.5
Vertex Cover
14
30
6.42,1 3.42 (# of nodes, # of edges)
12.6, 40.47
NP-Complete
N-Queens
15
30
3.8$\times$3.8 (grid size)
6.33$\times$6.33
NP-Hard Gent et al. (2017)
Knapsack
15
30
4.8 (array size)
24.56
NP-Hard
K Metric Centre
15
30
4.5 (# of nodes)
7
NP-Hard
Table 5: Names of problems in the PuzzleBench, number of samples in the training set, number of samples in the test set, average size of input instances in training set, average size of input instances in test set and computational complexity. The brackets in the 4th column describe how input instance sizes are measured. ? in the computational complexity column indicates that results are not available for the corresponding problem.

### v3

@@ 8 Discussion
1) Problems that SymPro-LM solved with 100% accuracy without any feedback. 8 such problems exist out of the $40$, including vertex-cover and latin-square. These problems have a one-to-one correspondence between the natural language description of the rules and the program for generating the constraints and the LLM essentially has to perform a pure translation task which they excel at.
@@
2) Problems that SymPro-LM solved with 100% accuracy but after feedback from solved examples. There are 20 such problems. They typically do not have a one-to-one correspondence between rule descriptions and code, thus requiring some reasoning to encode the problem in the solver’s language. For eg. one must define auxiliary variables and/or compose several primitives to encode a single natural language rule. GPT-4-Turbo initially misses constraints or encodes the problem incorrectly, but with feedback, it can spot its mistakes and corrects its programs. Examples include k-clique and binairo. In binairo, for example, GPT-4-Turbo incorrectly encodes the constraints for ensuring all columns and rows to be distinct but fixes this mistake after feedback (see Figure 17 in the appendix). LLMs can leverage their vast pre-training to discover non-trivial encodings for several interesting problems and solved examples can help guide LLMs to correct solutions in case of mistakes.
@@
3) Problems with performance below 100% that are not corrected through feedback or utilizing multiple runs. For these 12 problems, LLM finds it difficult to encode some natural language constraint into SMT. Examples include number-link and hamiltonian path, where GPT-4-Turbo is not able to figure out how to encode existence of paths as SMT constraints. In our opinion, these conversions are peculiar, and may be hard even for average CS students. We hope that further analysis of these 12 domains opens up research directions for neuro-symbolic reasoning with LLMs.
9 Conclusion and Limitations
@@
We investigate the reasoning abilities of LLMs on structured first-order combinatorial reasoning problems. We formally define the task, and we present FCoReBench, a novel benchmark of $40$ such problems and find that existing tool-augmented techniques, such as Logic-LM and PAL fare poorly. In response, we propose SymPro-LM – a new technique to aid LLMs with both program interpreters and symbolic solvers. It uses LLMs to convert text into executable code, which is then processed by interpreters to define constraints, allowing symbolic solvers to efficiently tackle the reasoning tasks. Our extensive experiments show that SymPro-LM’s integrated approach leads to superior performance on our dataset as well as existing benchmarks. Error analysis reveals that SymPro-LM struggles for a certain class of problems where conversion to symbolic representation is not straightforward. In such cases simple feedback strategies do not improve reasoning; exploring methods to alleviate such problems is a promising direction for future work. Another future work direction is to extend this dataset to include images of inputs and outputs, instead of serialized text representations, and assess the reasoning abilities of vision-language models, like GPT4-V.
Limitations: While we study a wide variety of fcore problems, more such problems always exist and adding these to FCoReBench remains a direction of future work. Additionally we assume that input instances and their outputs have a fixed pre-defined (serialized) representation, which may not always be easy to find. Another limitation is that encoding of many problems in the solver’s language can potentially be complicated. Our method relies on the pre-training of LLMs to achieve this without any training/fine-tuning, and addressing this is a direction for future work.
Appendix A FCoReBench
@@ A.1 Dataset Details and Statistics
Our dataset namely FCoReBench has $40$ different fcore problems that have been collected from various sources. Some of these problems are logical-puzzles from publishing houses like Nikoli, some problems are from operations research literature, some are from the annual SAT competition and other problems are well-known computational problems from Computer Science literature such as hamiltonian path and minimum-dominating set. Table 7 gives the details of all problems in our dataset. To create our training and test sets, we write scripts to synthetically generate problem instances. These can be used to extend the dataset as needed with any number of instances of any size. For experimentation, we generate some solved training instances and a separate set of testing instances. Each problem also has a natural language description of its rules, and a natural language description of the input-format which specify how input problem instances and their solutions are represented in text. The next few sections give illustrative examples and other details.
@@ A.2 Natural Language Description of Rules
This section describes how we create the natural language description of rules for problems in FCoReBench. We extract rules from the sources such as the Wikipedia/Nikoli pages of the corresponding problems. These rules are reworded by a human expert to reduce dataset contamination. Another human expert ensures that there are no ambiguities in the reworded description of the rules. The rules are generalized, when needed (for eg. from a $9\times 9$ Sudoku to a $n\times n$ Sudoku). The following sections provide few examples.
@@ A.2.1 Example Problem: Survo
Survo (Figure 7) is an example problem from FCoReBench. The task is to fill a $m\times n$ rectangular board with numbers from $1-m*n$ such that each row and column sums to an intended target. (Survo-Wikipedia). The box given below describes the rules of Survo more formally in natural language. {mdframed} We are given a partially filled $m\times n$ rectangular board, intended row sums and column sums.
- Empty cells are to be filled with numbers
- Numbers in the solved board can range from $1$ to $m*n$
- Numbers present in filled cells on the input board cannot be removed
- Each number from 1 to m*n must appear exactly once on the solved board
- All the empty cells should be filled such that each row and each column of the solved board must sum to the respective row sum and column sum as specified in the input
@@ A.2.2 Example Problem: Hamiltonian Path
{mdframed}
We are given an un-directed and un-weighted graph.
- We have to determine if the graph contains a path that visits every vertex exactly once.
@@ A.2.3 Example Problem: Dosun Fuwari
Dosun Fuwari (Nikoli) as shown in Figure 9 is another example problem from FCoReBench. We are given a square board with regions (cells enclosed in bold lines) and we have to fill the board with balloons and iron balls such that one balloon and one iron ball is placed in each region. Balloons are light and float, so they must be placed in one of the cells at the top, in a cell right under a black cell (filled-in cell), or under other balloons. Iron balls are heavy and sink, so they must be placed in one of the cells at the bottom, or in a cell right over a black cell or over other iron balls. The box given below gives the more formal description of the rules of dosun fuwari in natural language.
{mdframed}
We are given a partially filled n*n square board. We are also given subgrids of the input board. Cells in the input board can either be empty or filled (that is, nothing can be placed in them, they are blackened) or can be balloons or iron balls.
- The only thing we can do is place balloons or iron balls in some of or all of the empty cells
- Each subgrid specified in the input should have exactly one balloon and iron ball in the solved board
- Because balloons are buoyant, they should be positioned either in one of the cells located at the top of the board or in a cell directly below a filled cell (i.e., one of the blackened cells in the input) or below other balloons.
- Iron balls, being dense, will sink and should therefore be positioned either directly on one of the cells located at the bottom of the input board, or on a cell directly above a filled cell (i.e., one of the blackened cells in the input), or above another iron ball.
@@ A.3 Natural Language Description of Input and Output format
For many problems we consider input-output instances are typically not represented in text. For each problem we describe a straightforward conversion of the input and output space to text in natural language. The following sections consider examples of a few problems from FCoReBench.
@@ A.3.1 Example Problem: Survo
{mdframed}
Input Format:
- The input will have $m+1$ lines
- The first $m$ lines will have $n+1$ space-separated integers
- Each of these $m$ lines represents one row of the partially solved input board (n integers), followed by the required row sum (a single integer)
- The last line of the input will have $n$ space-separated integers each of which represents the required column sum in the solved board
Sample Input:
0 6 0 0 0 30
8 1 0 0 0 17
0 9 3 0 30
27 16 10 25
Output Format:
- The output should have $m$ lines, each representing one row of the solved board
- Each of these $m$ lines should have $n$ space-separated integers representing the cells of the solved board
- Each integer should be from $1$ to $m*n$
Sample Output:
12 6 2 10
8 1 5 4
7 9 3 11
@@ A.3.2 Example Problem: Dosun Fuwari
Figure 11 represents conversion of the inputs to dosun fuwari, originally represented as grid images to text. Here the first few lines represent the input board followed by a string ’—–’ which acts as a separator following which each of the lines has space-separated integers representing the subgrids of the input board. Cells are numbered in row-major order starting from 0, and this numbering is used to represent cells in each of the lines describing the subgrids. In the first few lines representing the input board, 0’s represent the empty cells that must be filled. 1’s denote the blackened cell, 2s denote the balloons and 3’s denote the iron balls. The box below describes these rules more formally in natural language {mdframed} Input-Format:
- The first few lines represent the input board, followed by a line containing --------, which acts as a separator, followed by several lines where each line represents one subgrid
- Each of the lines representing the input board will have space-separated integers ranging from 0 to 3
- 0 denotes empty cells, 1 denotes a filled cell (blackened cell), 2 denotes a cell with a balloon, 3 denotes a cell with an iron ball
- After the board, there is a separator line containing --------
- Each of the following lines has space-separated elements representing the subgrids on the input board
- Each of these lines has integers representing cells of a subgrid
- Cells are numbered in row-major order starting from 0, and this numbering is used to represent cells in each of the lines describing the subgrids
Sample-Input:
0 0 0 0
0 1 0 0
0 0 0 0
0 0 0 1
0 1
2 3 6 10
4 8 12
5 9 13 14 15
7 11
Output Format:
- The output should contain as many lines as the size of the input board, each representing one row of the solved board
- Each row should have n space separate integers (ranging from 0-3) where n is the size of the input board
- Empty cells will be denoted by 0s, filled cells (blackened) by 1s, balloons by 2s and iron balls by 3s
Sample-Output:
2 3 0 2
2 1 0 2
0 2 3 3
3 0 3 1
@@ A.3.3 Example Problem: Hamiltonian Path
Figure 12 represents the conversion of inputs to hamiltonian-path, originally represented as graph image to text. The first line denotes the number of vertices present in the graph followed by which each node of the graph will be numbered from 0 - N-1. Each of the subsequent lines represents an edge of the graph and will contain two space-separated integers (according to the numbering defined previously). The output is a single word (YES/NO) indicating if a hamiltonian path exists in the graph. The box below describes this more formally in natural language. {mdframed} Input Format:
- The first line will contain a single integer N, the number of nodes in the graph
- The nodes of the graph will be numbered from 0 to N-1
- Each of the subsequent lines will represent an edge of the graph and will contain two space-separated integers (according to the numbering defined above)
Sample-Input:
5
0 1
1 2
2 3
3 4
Output Format:
- The output should contain a single line with a single word
- The word should be YES if a path exists in the input graph according to constraints specified above and NO otherwise
Sample Output:
YES
Table 7: Names of problems in FCoReBench, number of samples in the training set, number of samples in the test set, average size of input instances in training set, average size of input instances in test set and computational complexity. The brackets in the 4th column describe how input instance sizes are measured. ? in the computational complexity column indicates that results are not available for the corresponding problem.
@@
3-Partition (Non Decision) 15 30 12 (array size) 17.7 NP-Hard
3-Partition (Decision) 15 30 12 (array size) 17.7 NP-Complete
Binario 15 50 4.0$\times$4.0 (grid size) 6.96$\times$6.96 NP-Hard [De Biasi, 2013]
Car-Sequencing 15 30 6.96, 3.66, 4.33 (# of cars, # of options, # of classes) 9.06, 5.66, 6.33 NP-Hard [Kis, 2004]
Clique Cover 15 30 6.26, 9.4 (# of nodes, # of edges) 12.9, 31.4 NP-Complete
Cryptarithmetic 15 30 4.32 (Average # of digits in the two operands ) 4.26 NP-Hard [Epstein, 1987]
Dosun Fuwari 15 30 3.066$\times$3.066 (grid size) 5.23$\times$5.23 NP-Hard [Iwamoto and Ibusuki, 2018]
Futoshiki 15 47 5$\times$5 (grid size) 7.57$\times$7.57 NP-Hard [Lloyd et al., 2022]
Fill-a-pix 15 35 $2.87\times 2.87$ (grid size) $4.1\times 4.1$ NP-Hard [HIGUCHI and KIMURA, 2019]
Flow-Shop 15 30 6.06, 3.4 (# of jobs, #num of machines) 3.83, 9.13 NP-Complete [Garey et al., 1976a]
Factory Workers 15 30 5.73, 12.66 (# of factories, # of workers) 12.35, 30.0 ?
Graph Coloring 15 30 5.13, 6.8 (# of nodes, # of edges) 9, 21.06 NP-Complete [Gent et al., 2017]
Hamiltonian Path 15 30 5.93, 8.6 (# of nodes, # of edges) 13.0, 19.77 NP-Complete
Hamiltonian Cycle 15 30 5.93, 8.6 (# of nodes, # of edges) 11.07, 18.67 NP-Complete
Hidato 15 45 $2.87\times 2.87$ (grid size) $4.1\times 4.1$ NP-Hard [Itai et al., 1982]
Independent Set 12 30 5.8, 7.2 (# of nodes, # of edges) 14.2, 29.8 NP-Complete
Inshi-No-Heya 15 49 5.0$\times$5.0 (grid size) 6.5$\times$6.5 ?
Job-Shop 15 30 3.66, 3.66 (# of jobs, # of machines) 9, 9 NP-Complete [Garey et al., 1976b]
K-Clique 15 31 4.87, 7.6 (# of nodes, # of edges) 8.84, 26.97 NP-Complete
Keisuke 15 30 4.33$\times$4.33 (grid size) 5.83$\times$5.83 ?
Ken Ken 15 20 3.26$\times$3.26 (grid size) 5.2$\times$5.2 NP-Hard [Haraguchi and Ono, 2015]
Knapsack 15 30 4.8 (array size) 24.56 NP-Hard
K Metric Centre 15 30 4.5 (# of nodes) 7 NP-Hard
Latin Square 15 50 6$\times$6.0 (grid size) 14.3$\times$14.3 NP-Hard [Colbourn, 1984]
Longest Path Problem 15 30 6.2, 5.87 (# of nodes, # of edges) 12.6, 16.3 NP-Complete
Magic Square 15 30 3.0$\times$3.0 (grid size) 4.$3$
Minimum Dominating Set 15 30 6.0, 17.73 (# of nodes, # of edges) 14.53, 45.0 NP-Complete
N-Queens 15 30 3.8$\times$3.8 (grid size) 6.33$\times$6.33 NP-Hard [Gent et al., 2017]
Number Link 15 50 4$\times$4 (grid size) 7.1$\times$7.1 NP-Hard
Partition Problem 15 35 7.06 (array size) 15 NP-Complete
PRP 15 30 4.93, 12.6 (# of units, # of days) 6.7, 23.9 ?
Shinro 15 30 5.13$\times$5.13 (grid size) 9.2$\times$9.2 ?
Subset Sum 15 30 3.67 (array size) 11.87 NP-Complete
Summle 15 20 2.33 (# of equations) 3.75 ?
Sudoku 15 50 4.0$\times$4.0 (grid size) 13.3$\times$13.3 NP-Hard [YATO and SETA, 2003]
Sujiko 15 45 3.0$\times$3.0 (grid size) 4.0$\times$4.0 ?
Survo 15 47 13.5 (area of grid) 20.25 ?
Symmetric Sudoku 15 30 4$\times$4 (grid size) 6.5$\times$6.5 ?
Sliding Tiles 15 30 $2.66\times 2.66$, 6.13 (grid size, search depth) $3.63\times 3.63$, 8.83 NP-Complete [Demaine and Rudoy, 2018]
Vertex Cover 14 30 6.4, 13.4 (# of nodes, # of edges) 12.6, 40.4 NP-Complete